# The DSA Transparency Database: Auditing Self-reported Moderation Actions by Social Media

### v2

@@ 5.2. RQ2: Data Inconsistencies
We found several inconsistencies in the self-reporting of social media platforms, based on a comparison between their DSA Transparency Reports and their SoRs. According to our comparative analysis, Pinterest resulted to be the most consistent in this regard, while also offering detailed data on its moderation actions. YouTube was also consistent in its self-reporting data, albeit the prominent spikes of restricted ad content in early October and mid December 2023 cast doubts on the completeness of the submitted data. The peculiar inconsistency of Snapchat, which indicated hundreds of notifications of illegal content by Trusted Flaggers when none has been yet designated under the DSA, seems to be an honest mistake due its internal namesake program. Nonetheless this peculiarity stands out as an example of the complexity that emerges when adapting already established internal processes to new and vast norms.
The majority of inconsistencies that we found was related to the use of automated means. In some cases, such inconsistencies appear to be a matter of interpretation. For instance, Facebook states in its transparency report: “[e]very day, we remove millions of violating pieces of content and accounts on Facebook. In most cases, this happens automatically, with technology to detect, restrict, and remove content and accounts […]”. Later, they further specify: “when confident enough that a post violates one of our Community Standards, the artificial intelligence will typically remove the content or demote it. We also use artificial intelligence to select the content for human review on the basis of severity, virality, and likelihood of a violation”. Instagram’s report includes similar statements. In other words, both Meta platforms declared heavy use of fully automated means in their transparency reports, but not a single fully automated action is reported in their submissions to the DSA-TDB, as shown in Figure 8. Instead, they opted to indicate partially automated detection by an overwhelming majority. Perhaps this choice was based on Meta’s declarations of “us[ing] human reviewers to assess the accuracy [of automated means] against current content, behaviour, or accounts, rather than just historical ones”. Regardless of the reason, we beg to differ, as we believe that at least some SoRs should have been classified as fully automated, because —as implied by Meta itself— there is no human involvement during the moderation action itself.
@@
Instead, X emerges as a distinctive outlier and as the most prominent case in self-reported inconsistencies. In its transparency report, the platform states: “X employs a combination of heuristics and machine learning algorithms to automatically detect content that violates the X Rules and policies […]”. Later they reiterate: “[c]ontent flagged by these machine learning models are either reviewed by human content reviewers before an action is taken or, in some cases, automatically actioned based on model output”. Yet, all of the SoR submissions by X during the period of study specified non-automated detection and decision. Not even a single partially automated decision was reported, as shown in Figure 8, which represents a striking inconsistency. In addition to the lack of use of automated means, other discrepancies include a lower than expected number of moderation actions, the unlikely predominance of illegal content, as well as a single value for both content type and restriction decision. These discrepancies in number and characterization cast serious doubts on the reliability of the SoRs submitted by X. These concerns appear to be well-founded, considering that during the same period in which we conducted our analysis, the European Commission initiated formal proceedings to evaluate whether X may have violated the DSA in aspects including, among others, content moderation practices (Commission, 2023a). On the other hand, X’s transparency report is arguably one of the most, if not the most, comprehensive of the eight platforms, with detailed tables and explanations regarding their moderation actions.
@@ 5.4. Limitations
Given the newness and complexities of the DSA in general and the DSA-TDB in particular, our study presents important limitations. One is the lag of several months between the last available transparency reports we used and the 100-day period of study. Even in this relatively short time frame, it is possible that the user share of platforms or their moderation practices have changed noticeably, explaining in part some of the inconsistencies we reported. However, we believe that if that were the case, our overall findings would still hold. In addition, this limitation is intrinsic to the nature of the reports themselves, which are expected to be published a few months after the end of their reference period. Another limitation related to the transparency reports is the lack of a standard format, which renders a direct comparison among them and with the data in the DSA-TDB particularly challenging, and even impossible in certain aspects due a lack of detail in some reports. In this regard, we think that DSA Transparency Reports should follow a similar structure, and at least use the common terminology already established for the DSA-TDB. Another related important limitation concerns the lack of platforms’ data outside the DSA-TDB to conduct data validation or more advanced analyses. For instance, and as already mentioned, we used MAAR as a proxy metric to relativize the volume SoRs, but it would have been more useful to also utilize the volume of pieces by content type created in the same time frame, which is not released by the eight analyzed platforms. Furthermore, we initially intended to conduct more analyses on the SoR data, but as stated before, we realized upon exploration that the records were not filled in as expected, in addition to the several inconsistencies therein. Hence, we refrained from conducting additional analysis to avoid obtaining potentially misleading results.
5.5. Final Remark: A Word of Caution
@@
Based on our results, we conclude that great care is needed when conducting research studies that draw upon the initial data of the DSA-TDB. Above all, one must not forget that these data are self-reported. Even if SoR submission and transparency reporting are mandatory under the DSA, so far the reliability of the data lies on the willingness and the resources of the platforms to be comprehensive and accurate in their mapping of existing moderation processes to the novel and complex norms. As we have shown, this is not yet the case for the nascent DSA-TDB. Indeed, our analyses indicate that social media SoRs are only partially adequate with respect to the objectives of the database, manifest critical inconsistencies, and reflect platforms with very different scopes, domains, and dynamics. Consequently, there is a high risk of making misleading inferences based on these data if the appropriate precautions are not taken, such as using complementary data sources for validation. Moreover, it is possible that these issues will persist or even be exacerbated once all providers of online platforms subject to the DSA —and not only VLOPs and VLOSEs— will be obliged to submit their moderation data to the DSA-TDB (European Parliament and Council, 2022). Considering that large, resourceful, and supposedly well-organized platforms experienced the many issues discussed herein, we expect that smaller and less resourceful platforms will struggle to do any better.
@@ Acknowledgements.
This work is partially supported by the European Union – Next Generation EU within the PRIN 2022 framework project PIANO (Personalized Interventions Against Online Toxicity).
@@ Appendix A Sources of DSA Transparency Reports
Platform name URL of document or repository

### v4

@@ 5.2. RQ2: Data Inconsistencies
We found several inconsistencies in the self-reporting of social media platforms, based on a comparison between their DSA Transparency Reports and their SoRs. According to our comparative analysis, Pinterest resulted to be the most consistent in this regard, while also offering detailed data on its moderation actions. YouTube was also consistent in its self-reporting data, albeit the prominent spikes of restricted ad content in early October and mid December 2023 cast doubts on the completeness of the submitted data. The peculiar inconsistency of Snapchat, which indicated hundreds of notifications of illegal content by Trusted Flaggers (§7.1 Information source) when none had been designated under the DSA at the time, seems to be an honest mistake due its internal namesake program. Nonetheless this peculiarity stands out as an example of the complexity that emerges when adapting already established internal processes to new and vast norms.
The majority of inconsistencies that we found were related to the use of automated means (§9 Automated detection and §10 Automated decision). In some cases, such inconsistencies appear to be a matter of interpretation. For instance, Facebook states in its transparency report: “[e]very day, we remove millions of violating pieces of content and accounts on Facebook. In most cases, this happens automatically, with technology to detect, restrict, and remove content and accounts […]”. Later, they further specify: “when confident enough that a post violates one of our Community Standards, the artificial intelligence will typically remove the content or demote it. We also use artificial intelligence to select the content for human review on the basis of severity, virality, and likelihood of a violation”. Instagram’s report includes similar statements. In other words, both Meta platforms declared heavy use of fully automated means in their transparency reports, but not a single fully automated action is reported in their submissions to the DSA-TDB, as shown in Figure 8. Instead, they opted to indicate partially automated detection by an overwhelming majority. Perhaps this choice was based on Meta’s declarations of “us[ing] human reviewers to assess the accuracy [of automated means] against current content, behaviour, or accounts, rather than just historical ones”. Regardless of the reason, we beg to differ, as we believe that at least some SoRs should have been classified as fully automated, because —as implied by Meta itself— there is no human involvement during the moderation action itself.
@@
Instead, X emerges as a distinctive outlier and as the most prominent case of self-reported inconsistencies. In its transparency report, the platform states: “X employs a combination of heuristics and machine learning algorithms to automatically detect content that violates the X Rules and policies […]”. Later they reiterate: “[c]ontent flagged by these machine learning models are either reviewed by human content reviewers before an action is taken or, in some cases, automatically actioned based on model output”. Yet, all of the SoR submissions by X during the period of study specified non-automated detection and decision. Not even a single partially automated decision was reported, as shown in Figure 8, which represents a striking inconsistency. In addition to the lack of use of automated means, other discrepancies include a lower than expected number of moderation actions, the unlikely predominance of illegal content, as well as a single value for both content type and restriction decision. These discrepancies in number and characterization cast serious doubts on the reliability of the SoRs submitted by X. These concerns appear to be well-founded, considering that during the same period in which we conducted our analysis, the European Commission initiated formal proceedings to evaluate whether X may have violated the DSA in aspects including, among others, content moderation practices (Commission, 2023a). On the other hand, X’s transparency report is arguably one of the most, if not the most, comprehensive of the eight platforms, with detailed tables and explanations regarding their moderation actions.
[…]
Square panel in which the centroid of the platforms’ characteristics is almost at the center. X is the platform furthest from the centroid, The second furthest is TikTok, but in a different quadrant. Pinterest is the third most distant platform, in yet another quadrant. Facebook and Instagram are close to one another but somewhat distant from the centroid. LinkedIn and YouTube are closer to the centroid and also somewhat close to one another. Finally, Snapchat is the closest platform to the centroid.
@@ 5.3. RQ3: Platform Differences
5.4. Suggestions to Improve the DSA Transparency Database
Despite the leap forward in terms of public accountability of online platforms brought by the DSA-TDB, its initial schema suffers some significant issues that hinder its usefulness for analysis and research. As we have mentioned before, we are aware of the enormous challenge that is to develop a single database that covers the various moderation practices of many heterogeneous online platforms at a continental scale. Based on our analyses, we have identified four issues that we believe should be fixed in future iterations of the database schema. To tackle these, we make the suggestions detailed in the following paragraphs, taking into account current practices of online platforms, the data inconsistencies that we found, the contents of the DSA itself, as well as the DSA-TDB documentation.
5.4.1. Make more attributes for decision grounds mandatory
As we previously detailed in Section 5.1, there are several attributes that can be used to describe the grounds for decision concerning a SoR, but for the most part these are underutilized, particularly the optional category addition and category specification. Hence, we believe that the DSA-TDB should at least make mandatory the currently optional and multi-valued category specification. This relatively small change to the database schema would oblige platforms to provide more context regarding their SoRs and it would greatly ease the analysis of grounds for decision, especially for values in category that have wide-ranging descriptions, such as scope of platform service.
5.4.2. Differentiate among cause and target entities of moderation decisions
In practice, moderation actions by online platforms usually involve three different entities: a single piece of content, an account, or a community group. We could say that a moderation decision is applied to a target entity due to the infringement that an associated cause entity represents. In most instances, the cause and target entities of a decision are the same. Oftentimes, however, there is an overlap among these, such as when the removal of an account or community group is also accompanied by the removal of their contents. In the most common case, platforms apply a moderation decision to a single piece of content because of the piece itself. Other times, if the piece of content is deemed a severe violation, the decision is applied to the account in addition to the infringing content. Both of these cases are mostly covered by the current structure of the DSA-DTB. Nonetheless, sometimes a decision is applied to an account by its behavior and not due to a specific piece of content, as in the case of repeated violations and the impersonation of people and companies not admissible under parody laws. In addition, a moderation decision might be applied to a community group without directly affecting the accounts that participated therein. These other cases are not well covered by the current DSA-TDB, forcing platforms to awkwardly map them to the available attributes. For instance, as illustrated in Figure 4, most of Snapchat’s SoRs with content type set to other are accounts. Similarly, for Facebook and Instagram all other content types are accounts, which represent more than half of their total SoRs. Moreover, Article 17(1) of the DSA states that SoRs must be provided by platforms “on the ground that the information provided by the recipient of the service is illegal content or incompatible with their terms and conditions”. We argue that the term information goes beyond the mere content created by the recipient —e.g., it could be false information provided by the recipient when creating an impersonation account. Therefore, in spite of the relatively important impact on the database schema that such a differentiation among these entities represents, we believe that this change is justified, especially for what concerns differentiating between content and account.
5.4.3. Flag synthetic content and accounts separately
We argue that the current use of the value synthetic media for the attribute content type is semantically inappropriate and prone to overuse by online platforms, as was the case with X (see Figure 4). In addition, accounts deemed to be synthetic might mainly interact with non-synthetic content, such as via content engagement botnets for hire, for which specifying synthetic content on the respective SoR might not be suitable. Instead, we think that the cause entity of a moderation decision (e.g., content, account) should have a separate boolean attribute synthetic. Thanks to this relatively small change, it would be possible to better analyze the multimedia types that are used in contents and accounts deemed synthetic.
5.4.4. Clarify and expand attributes for automated means
We think that the two attributes related to the use of automated means (detection and decision) are not well specified, which gave way to different interpretations by the analyzed platforms, as we detailed in Sections 4.5 and 4.6. For instance, the official DSA-TDB auxiliary documentation titled Additional Explanation For Statement Attributes [^12] https://transparency.dsa.ec.europa.eu/page/additional-explanation-for-statement-attributes reads in § 9 Automated detection: “This attribute automated_detection indicates whether and to what extent automated means were used to identify the specific information addressed by the decision. ‘Yes’ means that automated means were used to identify the specific information addressed by the decision”. We argue that the current dichotomous attribute does not actually indicate to what extent automated means were used to identify infringing information. It merely indicates their use or not. Instead, we suggest using an ordinal attribute with the same values as automated decision: fully automated, not automated, and partially automated. We also suggest a more detailed definition of the previous three values. In § 10 Automated detection of the aforementioned auxiliary documentation, it is stated that “‘Fully automated’ means that the entire decision-process was carried out without human intervention.” We did not find, however, a definition of the decision-process to which this statement refers within the DSA-TDB documentation, and we believe this to be the main source of variation in interpretation of automated means among platforms and even within the same platform. This seems to be the case of both Meta’s platforms, which declare a heavy use of fully automated decisions in their transparency reports but afterward declared none to the DSA-TDB, opting instead to mostly submit SoRs as partially automated decisions. We believe that the above decision-process should be delimited by interaction with the infringing entity of the SoR itself. If a human actor was not directly involved in the detection (e.g., by flagging) or decision (e.g., by reviewing) applied to the specific piece of content, account, or community group associated with the given SoR in which automated models or algorithms were used, then the respective means should be considered fully automated. This small change in schema and documentation would provide a consistent and more nuanced scale across both automated means. It would also reduce the current ambiguity regarding the related yet separate prior development (and refinement) process of the automated means themselves (e.g., algorithmic specification or training of AI models), which in most, if not all, cases involves a human actor at some point or another.
@@ 5.5. Limitations
Given the newness and complexities of the DSA in general and the DSA-TDB in particular, our study presents important limitations. One is the lag of several months between the last available transparency reports we used and the 100-day period of study. Even in this relatively short time frame, it is possible that the user share of platforms or their moderation practices have changed noticeably, explaining in part some of the inconsistencies we reported. However, we believe that if that were the case, our overall findings would still hold. In addition, this limitation is intrinsic to the nature of the reports themselves, which are expected to be published a few months after the end of their reference period. Another limitation related to the transparency reports is the lack of a standard format, which renders a direct comparison among them and with the data in the DSA-TDB particularly challenging, and even impossible in certain aspects due a lack of detail in some reports. In this regard, we think that DSA Transparency Reports should follow a similar structure, and at least use the common terminology already established for the DSA-TDB. Another related important limitation concerns the lack of platforms’ data outside the DSA-TDB to conduct data validation or more advanced analyses. For instance, and as already mentioned, we used MAAR as a proxy metric to relativize the volume SoRs, but it would have been more useful to also utilize the volume of pieces by content type created in the same time frame, which is not released by the eight analyzed platforms. Furthermore, we initially intended to conduct more analyses on the SoR data, but as stated before, we realized upon exploration that the records were not filled in as expected, in addition to the several inconsistencies therein. Hence, we refrained from conducting additional analysis to avoid obtaining potentially misleading results. Among the fields that we did not consider, but that could be worth analyzing in future research, are the duration of the moderation actions (§4 The duration of the restriction), their territorial scope (§5 The territorial scope of the decision), and the free text description of the facts and circumstances that led to the moderation (§6 Description of the facts and circumstances).
5.6. Implications for Future Research on Online Content Moderation
The launch of the DSA-TDB marks a major milestone in regulation and accountability of online content moderation. Albeit it is still too early to confidently evaluate the impact of this database and the DSA as whole on industry, academia, and society in general, in the following paragraphs we ponder on some of the most important implications for research on online moderation, both in the medium and long terms.
5.6.1. Unprecedented Access to Large Scale Moderation Data
One of the key aspects of the DSA-TDB is the public availability of the full archive data, all for free and without restrictions. This means that anyone, including industry and scholarly researchers, could analyze these data to better understand the dynamics of online moderation in the EU. It should be noted, however, that the data on the DSA-TDB is intentionally not exhaustive, due to privacy concerns and the scope of the related Article 17 of the DSA, which only covers statements of reasons that platforms must provide to recipients when a restriction is applied to their information on the platform. The DSA provides, however, a different mechanism for vetted researchers to gain access to more comprehensive data directly from VLOPs and VLOSEs by means of Article 40(12), subject to an application and approval process. As part of this vetting process, interested researchers must meet several requirements concerning, among other, their affiliation, research scope on systemic risk (as defined in Article 34) and/or risk mitigation (as defined in Article 35), ability to fulfil data security and confidentiality, and commitment to publicly release the derived research outcomes free of charge. Notwithstanding the stringent requirements laid out in Article 40, such mechanism allows a more fine-grained and comprehensive scrutiny of the most prominent online platforms, and it could also be used to complement the data available on the TSA-DBA. At the moment of writing (mid 2024), the affected social media platforms have already put in place their vetting process. Nonetheless, several researchers have raised concerns regarding the full compliance of the implemented processes (Jaursch et al., 2024). Furthermore, the European Commission has initiated formal proceedings against the platform X, because —among other preliminary findings— it has so far failed to provide access to approved researchers as required by Article 40(12). [^13] https://ec.europa.eu/commission/presscorner/detail/en/IP_24_3761 The actual consequences for very large platforms that do not comply with the DSA and their effectiveness remain to be seen, however.
5.6.2. Common Conceptual Model for Online Moderation
It could be argued that the schema and documentation associated with the DSA-TDB codify a common core conceptual model to discuss online moderation across different kinds of platforms. Due to the large scope of the database, the definitions of the attributes and values therein are likely to be echoed across many derived scholarly works and beyond. Therefore, it is paramount that their description is clear and as unambiguous as possible for all the stakeholders involved in the production and consumption of these data. As we have shown in our analyses herein, this is not yet the case.
5.6.3. Risk of Over-representing European Social Media Practices
Several scholars have decried that despite their global reach, there is a prevalent US-centrism (and lesser Euro-centrism) in online social media platforms —in policies, content created, and studies performed (Nieborg et al., 2022). In fact, all of the eight platforms analyzed herein were founded and are headquartered in the United States. In view of the previous two implications regarding data availability and a common conceptual model via the DSA-TDB, it is very likely that in the following years there will be a remarkable increase in the share of online moderation studies based on data and policies from within the EU. Nonetheless, researchers and experts should refrain from generalizing the findings of these studies to other regions of the world, particularly low-income countries, which are usually under-represented in social media research but often consist of sizeable or even larger user bases, compared to most EU countries.
5.6.4. A Word of Caution on Initial Database Analyses
@@
Based on our results, we conclude that great care is needed when conducting research studies that draw upon the initial data of the DSA-TDB. Above all, one must not forget that these data are self-reported. Even if SoR submission and transparency reporting are mandatory under the DSA, so far the reliability of the data lies on the willingness and the resources of the platforms to be comprehensive and accurate in their mapping of existing moderation processes to the novel and complex norms. As we have shown, this is not yet the case for the nascent DSA-TDB. Indeed, our analyses indicate that social media SoRs are only partially adequate with respect to the objectives of the database, manifest critical inconsistencies, and reflect platforms with very different scopes, domains, and dynamics. Moreover, data inconsistency and lack of data in optional attributes makes it difficult to perform more robust quantitative analyses to derive solid conclusions or generalize findings. Consequently, there is a high risk of making misleading inferences based on these data if the appropriate precautions are not taken, such as using complementary data sources for validation. Furthermore, it is possible that the issues described herein persist or are even exacerbated now that all providers of online services subject to the DSA —and not only VLOPs and VLOSEs— are obliged to submit their moderation data to the DSA-TDB (European Parliament and Council, 2022). Considering that large, resourceful, and supposedly well-organized platforms experienced the many issues discussed herein, we expect that smaller and less resourceful platforms will struggle to do any better. The impact of Article 17 of the DSA on smaller platforms is however yet to be determined.
@@ Acknowledgements.
This work is partially supported by the European Union—NextGenerationEU within the ERC project DEDUCE (Data-driven and User-centered Content Moderation) under grant #101113826, and the PRIN 2022 project PIANO (Personalized Interventions Against Online Toxicity) under CUP B53D23013290006.
@@ Appendix A Sources of DSA Transparency Reports
platform URL of document or repository
@@ Appendix B Daily Count of Moderation Actions
[…]
Faceted time series. Facebook and Instagram reached their peaks round early December. Linkedin displays a weekly seasonality with a declining trend around the end of year holidays. Pinterest and Snapchat display a relatively stable trend over time. TikTok display two significant dips at the beginning of November and December, respectively. YouTube presents three remarkable spikes. The lowest at the beginning of October, and then the two most prominent in early December.